Consider nominal GDP is 1500 and the money supply is 400.

Instructions: Enter numerical values to two decimal places.

a) What is the velocity? V =

b) If nominal GDP rises to 1600, but the money supply does not change, how has velocity changed? Velocity will (increase/decrease) by to V =

c) If GDP now falls back to 1500 and the money supply falls to 350, what is velocity? V =
Answer #1

Answer

a)

the equation of exchange in the market is:

MV=PY

M=money supply

V=velocity

PY=nominal GDP

V=PY/V

V=1500/400=3.75

the velocity is 3.75

==========

b)

V=1600/400

=4

the velocity increases by =4-3.75=0.25

=========

c)

V=1500/350

=4.28571429

=4.29

the velocity is 4.29